# A JWST Transit of a Jupiter Analog I. Constraints on the Oblateness of Kepler-167 e

### v1

@@ Abstract
1 Introduction
Two decades ago, Hui & Seager (2002) concluded an early study on exoplanet transit modeling with the following passage: “…it is clear that future high-precision measurements of extrasolar planetary transits will present a very interesting challenge—there are several sources of small secondary fluctuations… that are within observational reach and that would require some effort to disentangle. The rewards of such efforts will be substantial—from detection of moons and rings to measuring the (projected) planet oblateness…”
@@ 1 Introduction
This article is the first in a series focused on the analysis of JWST observations collected in October of 2024 as a part of program GO:6491, “Revealing the Oblateness and Satellite System of an Extrasolar Jupiter Analog” (PI: Cassese). As described in Section 2, this remarkably-Jupiter like planet is one of the coldest, longest-period known transiting exoplanets, and also happens, fortunately, to transit a (relative for Kepler stars) bright, quiet, sub-Solar radius star. In other words, it is an ideal candidate to search for features that are common among the cold gas giants of our solar system but have yet to be observed beyond it, including Galilean-like moons or rotation-induced flattening. This article details our analysis concerning the planet’s oblateness, while Kipping et al. 2025 (hereafter Paper II) details our search for a Ganymede-like satellite.
To begin, we define planetary oblateness. In this context we refer to oblateness as the relative expansion of a planet’s equatorial radius relative to its polar radius caused by the centrifugal forces of the planet’s rotation (Barnes & Fortney, 2003). In this conceptualization, the planet’s equilibrium shape is an oblate spheroid, with two equal length axes in the equatorial planet $R_{\text{eq}}$ and a shorter polar radius $R_{\text{p}}$. We define the planet’s true oblateness $f^{*}$ as:
@@
In our solar system, Saturn is the most oblate of the gas giants with $f^{*}$ near 0.1, followed by Jupiter with $f^{*}$ near 0.05. These gas giants have rotation periods of 10.6 and 9.9 hours, respectively, which illustrates that rotation rate alone does not set a planet’s oblateness: instead, it instead also influenced by additional properties such as mass and internal structure. Barnes & Fortney (2003) show that by using the Darwin-Radau approximation to account for the redistribution of mass caused by the equatorial swelling, $f^{*}$ is related to the planet’s rotation by the following:
@@
Since we can generate approximations for $\mathbb{C}$ and constrain $R_{\text{eq}}$ and $M_{p}$ from measurements, Eq. 2 implies that we can map directly from $f^{*}$ to a planet’s rotation rate. This is a fundamental quantity capable of providing insights into giant planet formation and atmospheric dynamics that cannot be probed in any other way for planets not amenable to direct imaging (Batygin, 2018; Bryan et al., 2018). In addition to these rewards of measuring $f^{*}$, there are risks to not doing so as well. Namely, using a spherical planet model to infer the radius of an oblate planet will lead to a biased estimate of the planet’s radius, which in turn will bias inferences on the planet’s density and composition (Burton et al., 2014; Berardo & de Wit, 2022a).
@@
Several techniques have been proposed to probe $f^{*}$, including by comparing variations in transit depth between different epochs that may be caused by the precession of an oblate planet’s spin vector, which in turn would alter its projected area and resulting transit depth (Carter & Winn, 2010a; Biersteker & Schlichting, 2017). Similarly, precession could also shift the contact times of a transit, leading to a potentially detectably transit timing variation (Berardo & de Wit, 2022b). Alternatively, an oblate planet may imprint a detectable signal on a spectroscopic Rossiter–McLaughlin measurement (Akinsanmi et al., 2020).
@@
However, arguably the most straightforward way to probe $f^{*}$ is to search for deviations from spherical transit models on the ingress and egress of a transit. Since the moment of first contact between an oblate planet and the stellar limb will appear early or late relative to a spherical planet with the same projected radius, spherical planet models are unable to fully capture the complex shape of the ingress and egress (Barnes & Fortney, 2003) [^1] Note that once fully in transit, an oblate planet will appear nearly indistinguishable from a spherical planet of the same projected area. Nearly, but not exactly, since they’ll subtend slightly different portions of the star. Since the star’s intensity varies with projected radius (limb darkening), the total flux blocked by these different shapes will vary slightly. However, this effect is <1<1 ppm (Cassese et al., 2024). Therefore, the best-fitting spherical model will contain small (10s-100s of ppm, Cassese et al., 2024) “wiggles” in the residuals during ingress and egress. By forward modeling a more complex simulation of an oblate transistor, we can fit these wiggles, place constraints on the oblateness, and perform model comparison with a spherical planet model.
@@
There is an important caveat to this technique, however: it is only sensitive to $f$, oblateness of the on-sky projected ellipse, not $f^{*}$. This projected oblateness is set by the combination of the true oblateness and two rotation angles (the planet’s obliquity and rotation about the $\hat{z}$ axis of its orbital plane) and must necessarily be a lower bound on the true oblateness (Berardo & de Wit, 2022b). The mapping between true and projected oblateness is given in Carter & Winn (2010b), among others.
@@
Several studies have attempted to measure $f$ in the past, including Carter & Winn (2010b), who analyzed Spitzer data of HD 189733 b, Zhu et al. (2014), who analyzed a selection of Kepler planets, and (Price et al., 2025), who developed a new forward modeling package and applied it to K2 data of HIP-41378 f. Dholakia et al. (2025) recently created their own forward modeling package and applied it to JWST data of WASP-107 b and were able to place an upper bound of $f<0.23$. Finally Lammers & Winn (2024) used squishyplanet (Cassese et al., 2024), another open-source package for modeling the transits of non-spherical planets, to place an upper limit of $f<0.16$ for the super puff Kepler-51 d using JWST data as well. These latter three constraints were used to place limits on the rotation rate of $P\geq 15.3$, $P\geq 13$, and $P\geq 40$ hours for HIP-41378 f, WASP-107 b and Kepler-51 d, respectively. Note, however, that these limits are conditioned upon different assumptions about the true alignment of the planet relative to the sky frame. Dholakia et al. (2025) assume the projected and true oblateness are equal, while Lammers & Winn (2024) assumes that the planet’s true obliquity is $45^{\circ}$.
@@
We describe Kepler-167 e and our observations in Sec. 2. We detail the grid of reductions and trend models explored in Sec. 3. We then describe how we culled this extensive grid via maximum likelihood fits in Sec. 4. We describe our posterior inference via nested sampling across a smaller grid of models in Sec. 5. Finally, we offer our conclusions in Sec. 6.
2 Target and Observations
@@
Kepler-167 e which was first discovered by Kipping et al. (2016) through analysis of two transits observed by the Kepler mission. The planet has since been the target of three other observing campaigns prior to this study: Dalba & Tamburo (2019) recorded a partial transit with the Spitzer Space Telescope and managed to establish a workable ephemeris by ruling out large TTVs; next, a group of citizen scientists from nine different countries operating through the Unistellar Telescope Network recorded an additional transit (Perrocheau et al., 2022); and finally, Chachan et al. (2022) analyzed radial velocity measurements taken with Keck-HIRES from 2017-2020 to constrain the planet’s mass and eccentricity.
@@
Due to the extreme length of the time series and onboard processing/storage requirements of the observatory, our observations were split into 6 distinct exposures with a small pause between each. During this time, the detector was reset and briefly entered its passive continuous-readout mode. The JWST Engineering Mnemonic Database also indicates that the High Gain Antenna moved during these pauses, as well as several times within the science exposures. Exposures 1-2 contain no planetary transits; exposures 3-5 contain the transit of Kepler-167 e; exposure 6 contains a transit of Kepler-167 c, a super-Earth planet on a 7.4 day orbit (Rowe et al., 2014; Kipping et al., 2016). As seen in Fig. 1 and discussed in Sec. 3.4, each of these exposures were influenced by their own decreasing and nonlinear background trend. We note that no pixels are flagged as saturated by the jwst pipeline (Bushouse et al., 2025) in the median image.
@@
3 Initial Model Exploration
3.1 Summary of Approach
A planet’s oblateness, as discussed in Sec. 1, induces 10s-100s of ppm “wiggles” during a planet’s transit ingress and egress that cannot be captured by a spherical planet model (Barnes & Fortney, 2003). While at first brush the scale of this signal appears comparable with atmospheric features that JWST regularly targets (e.g., Alam et al., 2025), it is important to note a key difference between measuring oblateness and measuring variations in transit depth. While those chasing a planet’s spectrum can use the entire time series to inform their depth measurements, an oblateness signal presents itself on a much shorter timescale. The signatures of oblateness only imprint themselves on the light curve during ingress and egress, and even worse, are (generally) not symmetric across both. In the case of Kepler-167 e, the relevant timescale is $\sim 15-30$ min.
@@ 3.1 Summary of Approach
Our grid consisted of all combinations of the following: 3 different reduction pipelines, each ran with 6 different settings, for a total of 18 unique reductions; 6 unique binning cadences per reduction, producing a total of 108 unique light curves; 3 different heuristic trend models to capture exposure-long drifts; 3 different Gaussian Process additions to the trend model [^3] One of the trend options was simply a constant trend with the intention of lettering the GP handle all drifts, and one of the GP options was to leave off the GP. Consequently, we did not test models that used a constant trend but no GP.; and, finally, 5 different limb darkening models, including two novel parameterizations. All told we attempted to fit 4320 unique light curve/forward model combinations. To ensure our final likelihood values robustly reflect the true global minimum, we repeated each fit 3 times with 3 different initial random seeds. All fits were run on Yale University’s Bouchet high-performance computing cluster.
The remainder of this section is dedicated to a description of each of these models, while the results of the point-estimate fits are presented in Sec. 4.
3.2 Reduction Pipelines and Light Curve Generation
@@ 3.2.2 ExoTiC-JEDI
katahdin only includes functionality to process the equivalent of Stage 1 of jwst, meaning it includes no new algorithms for spectral extraction. In an effort to maintain true independence from the other pipelines, however, we opted not to use either jwst, Eureka!, or ExoTiC-JEDI for the equivalent of Stage 3 processing. Instead, we used the following simple routine: looking at just exposure 1, which does not contain any of Kepler-167 e’s transit, we flagged all pixels that fell above various signal-to-noise thresholds and summed them column-wise. This was justified by the minimal background signal left following the GP $1/f$ correction. Our 6 katahdin light curves correspond to using SNR thresholds 3, 5, 8, 10, 15, and 25, where as ordered each light curve relied on fewer and fewer pixels. As discussed in Sec. 4.4, however, we only used the latter four of these in our final nested sampling fits.
@@
3.2.4 Binning/Cleaning Procedure
@@
3.2.5 Pipeline Comparisons
Here we take the opportunity to discuss the white light curves themselves absent any reference to their influence on our fit results. Fig. 1 shows all 18 reductions plotted in full in their 5-minute cadence forms, each normalized by the median of flux in exposure 1. Fig. 2 shows a zoomed-in view of just exposure 1 after dividing out the best-fit quadratic trend from each time series. Two immediate features jump out here: first, light curves that were produced by the same pipeline tend to group together, regardless of their individual pipeline settings. Secondly, all three pipelines tend to agree with one another. Fig. 3 validates both of these conclusions by plotting the correlations between white light curves derived from the same pipeline vs. those from two different pipelines: both cases are positively correlated, but the inter-pipeline distributions are tighter.
@@ 3.2.5 Pipeline Comparisons
3.3 Limb Darkening Models
@@ 3.3.1 Uninformative Quadratic
3.3.2 Uninformative Cubic
Motivated by the shortcomings of quadratic limb darkening, we also included a more flexible model in our grid: cubic polynomials. Cubics present several practical challenges: for one, one of the most commonly used transit models, Mandel & Agol (2002), is explicitly built around quadratics. This is thankfully circumvented by using a modeling package such as squishyplanet that implements the arbitrary-order limb darkening framework presented in Agol et al. (2020).
@@ 3.3.2 Uninformative Cubic
3.3.3 Physically Informed Limb Darkening
@@
To do this, we start by downloading a subset of the pre-computed stellar atmosphere models that underpin exotic-ld (Grant & Wakeford, 2024), an open-source package for computing theoretical limb darkening profiles. Each file corresponds to a star with a certain $T_{\mathrm{eff}}$, log $g$, and metallicity; each file contains several spectra of that star corresponding to different distances from the projected center. We also download the appropriate instrument response curve for our specific filter/grating combination from exotic-ld.
@@
One may fairly ask in all of this why we do not simply fix the limb darkening coefficients before a fit as is occasionally done elsewhere in the literature (e.g. Alam et al., 2022). Doing so, however, runs the risk that a slight misparameterization could irreparably inject or suppress an oblateness signal (and/or alter a planet’s spectrum, Mercier et al., 2025). Further, even if one were to fully trust a given stellar atmosphere model (which would be unwise, given the evidence that they do not perfectly represent observed empirical profiles Espinoza & Jordán, 2015; Patel & Espinoza, 2022), every star comes with uncertainties on its physical characteristics as well.
@@
We have assembled this routine into an open-source python/JAX package we call asterias, after the Latin genus of the common star fish, and hope that the community finds it useful for modeling the limbs of stars. We emphasize that asterias is built on the work of Grant & Wakeford (2024) and their efforts to compile stellar atmosphere model outputs into standardized file types, and that any efforts that rely on asterias should also cite Grant & Wakeford (2024).

### v2

@@ Abstract
I Introduction
Two decades ago, Hui and Seager (2002) concluded an early study on exoplanet transit modeling with the following passage: “…it is clear that future high-precision measurements of extrasolar planetary transits will present a very interesting challenge—there are several sources of small secondary fluctuations… that are within observational reach and that would require some effort to disentangle. The rewards of such efforts will be substantial—from detection of moons and rings to measuring the (projected) planet oblateness…”
@@ I Introduction
This article is the first in a series focused on the analysis of JWST observations collected in October of 2024 as a part of program GO:6491, “Revealing the Oblateness and Satellite System of an Extrasolar Jupiter Analog” (PI: Cassese). As described in Section II, this remarkably-Jupiter like planet is one of the coldest, longest-period known transiting exoplanets, and also happens, fortunately, to transit a (relative for Kepler stars) bright, quiet, sub-Solar radius star. In other words, it is an ideal candidate to search for features that are common among the cold gas giants of our solar system but have yet to be observed beyond it, including Galilean-like moons or rotation-induced flattening. This article details our analysis concerning the planet’s oblateness, while Kipping et al. 2025 (hereafter Paper II) details our search for a Ganymede-like satellite.
To begin, we define planetary oblateness. In this context we refer to oblateness as the relative expansion of a planet’s equatorial radius relative to its polar radius caused by the centrifugal forces of the planet’s rotation (Barnes and Fortney, 2003). In this conceptualization, the planet’s equilibrium shape is an oblate spheroid, with two equal length axes in the equatorial planet $R_{\text{eq}}$ and a shorter polar radius $R_{\text{p}}$. We define the planet’s true oblateness $f^{*}$ as:
@@
In our solar system, Saturn is the most oblate of the gas giants with $f^{*}$ near 0.1, followed by Jupiter with $f^{*}$ near 0.05. These gas giants have rotation periods of 10.6 and 9.9 hours, respectively, which illustrates that rotation rate alone does not set a planet’s oblateness: instead, it instead also influenced by additional properties such as mass and internal structure. Barnes and Fortney (2003) show that by using the Darwin-Radau approximation to account for the redistribution of mass caused by the equatorial swelling, $f^{*}$ is related to the planet’s rotation by the following:
@@
Since we can generate approximations for $\mathbb{C}$ and constrain $R_{\text{eq}}$ and $M_{p}$ from measurements, Eq. 2 implies that we can map directly from $f^{*}$ to a planet’s rotation rate. This is a fundamental quantity capable of providing insights into giant planet formation and atmospheric dynamics that cannot be probed in any other way for planets not amenable to direct imaging (Batygin, 2018; Bryan et al., 2018). In addition to these rewards of measuring $f^{*}$, there are risks to not doing so as well. Namely, using a spherical planet model to infer the radius of an oblate planet will lead to a biased estimate of the planet’s radius, which in turn will bias inferences on the planet’s density and composition (Burton et al., 2014; Berardo and de Wit, 2022b).
@@
Several techniques have been proposed to probe $f^{*}$, including by comparing variations in transit depth between different epochs that may be caused by the precession of an oblate planet’s spin vector, which in turn would alter its projected area and resulting transit depth (Carter and Winn, 2010b; Biersteker and Schlichting, 2017). Similarly, precession could also shift the contact times of a transit, leading to a potentially detectably transit timing variation (Berardo and de Wit, 2022a). Alternatively, an oblate planet may imprint a detectable signal on a spectroscopic Rossiter–McLaughlin measurement (Akinsanmi et al., 2020).
@@
However, arguably the most straightforward way to probe $f^{*}$ is to search for deviations from spherical transit models on the ingress and egress of a transit. Since the moment of first contact between an oblate planet and the stellar limb will appear early or late relative to a spherical planet with the same projected radius, spherical planet models are unable to fully capture the complex shape of the ingress and egress (Barnes and Fortney, 2003) [^1] Note that once fully in transit, an oblate planet will appear nearly indistinguishable from a spherical planet of the same projected area. Nearly, but not exactly, since they’ll subtend slightly different portions of the star. Since the star’s intensity varies with projected radius (limb darkening), the total flux blocked by these different shapes will vary slightly. However, this effect is <1<1 ppm (Cassese et al., 2024). Therefore, the best-fitting spherical model will contain small (10s-100s of ppm, Cassese et al., 2024) “wiggles” in the residuals during ingress and egress. By forward modeling a more complex simulation of an oblate transistor, we can fit these wiggles, place constraints on the oblateness, and perform model comparison with a spherical planet model.
@@
There is an important caveat to this technique, however: it is only sensitive to $f$, oblateness of the on-sky projected ellipse, not $f^{*}$. This projected oblateness is set by the combination of the true oblateness and two rotation angles (the planet’s obliquity and rotation about the $\hat{z}$ axis of its orbital plane) and must necessarily be a lower bound on the true oblateness (Berardo and de Wit, 2022a). The mapping between true and projected oblateness is given in Carter and Winn (2010a), among others.
@@
Several studies have attempted to measure $f$ in the past, including Carter and Winn (2010a), who analyzed Spitzer data of HD 189733 b, Zhu et al. (2014), who analyzed a selection of Kepler planets, and (Price et al., 2025), who developed a new forward modeling package and applied it to K2 data of HIP-41378 f. Dholakia et al. (2025) recently created their own forward modeling package and applied it to JWST data of WASP-107 b and were able to place an upper bound of $f<0.23$. Finally Lammers and Winn (2024) used squishyplanet (Cassese et al., 2024), another open-source package for modeling the transits of non-spherical planets, to place an upper limit of $f<0.16$ for the super puff Kepler-51 d using JWST data as well. These latter three constraints were used to place limits on the rotation rate of $P\geq 15.3$, $P\geq 13$, and $P\geq 40$ hours for HIP-41378 f, WASP-107 b and Kepler-51 d, respectively. Note, however, that these limits are conditioned upon different assumptions about the true alignment of the planet relative to the sky frame. Dholakia et al. (2025) assume the projected and true oblateness are equal, while Lammers and Winn (2024) assumes that the planet’s true obliquity is $45^{\circ}$. Note also that Liu et al. (2024) analyzed the same JWST data underlying Lammers and Winn (2024) using their own forward model (Liu et al., 2025) and arrived at a more conservative bound of $f<0.2$.
@@
We describe Kepler-167 e and our observations in Sec. II. We detail the grid of reductions and trend models explored in Sec. III. We then describe how we culled this extensive grid via maximum likelihood fits in Sec. IV. We describe our posterior inference via nested sampling across a smaller grid of models in Sec. V. Finally, we offer our conclusions in Sec. VI.
II Target and Observations
@@
Kepler-167 e which was first discovered by Kipping et al. (2016) through analysis of two transits observed by the Kepler mission. The planet has since been the target of three other observing campaigns prior to this study: Dalba and Tamburo (2019) recorded a partial transit with the Spitzer Space Telescope and managed to establish a workable ephemeris by ruling out large TTVs; next, a group of citizen scientists from nine different countries operating through the Unistellar Telescope Network recorded an additional transit (Perrocheau et al., 2022); and finally, Chachan et al. (2022) analyzed radial velocity measurements taken with Keck-HIRES from 2017-2020 to constrain the planet’s mass and eccentricity.
@@
Due to the extreme length of the time series and onboard processing/storage requirements of the observatory, our observations were split into 6 distinct exposures with a small pause between each. During this time, the detector was reset and briefly entered its passive continuous-readout mode. The JWST Engineering Mnemonic Database also indicates that the High Gain Antenna moved during these pauses, as well as several times within the science exposures. Exposures 1-2 contain no planetary transits; exposures 3-5 contain the transit of Kepler-167 e; exposure 6 contains a transit of Kepler-167 c, a super-Earth planet on a 7.4 day orbit (Rowe et al., 2014; Kipping et al., 2016). As seen in Fig. 1 and discussed in Sec. III.4, each of these exposures were influenced by their own decreasing and nonlinear background trend. We note that no pixels are flagged as saturated by the jwst pipeline (Bushouse et al., 2025) in the median image.
@@
III Initial Model Exploration
III.1 Summary of Approach
A planet’s oblateness, as discussed in Sec. I, induces 10s-100s of ppm “wiggles” during a planet’s transit ingress and egress that cannot be captured by a spherical planet model (Barnes and Fortney, 2003). While at first brush the scale of this signal appears comparable with atmospheric features that JWST regularly targets (e.g., Alam et al., 2025), it is important to note a key difference between measuring oblateness and measuring variations in transit depth. While those chasing a planet’s spectrum can use the entire time series to inform their depth measurements, an oblateness signal presents itself on a much shorter timescale. The signatures of oblateness only imprint themselves on the light curve during ingress and egress, and even worse, are (generally) not symmetric across both. In the case of Kepler-167 e, the relevant timescale is $\sim 15-30$ min.
@@ III.1 Summary of Approach
Our grid consisted of all combinations of the following: 3 different reduction pipelines, each ran with 6 different settings, for a total of 18 unique reductions; 6 unique binning cadences per reduction, producing a total of 108 unique light curves; 3 different heuristic trend models to capture exposure-long drifts; 3 different Gaussian Process additions to the trend model [^3] One of the trend options was simply a constant trend with the intention of letting the GP handle all drifts, and one of the GP options was to leave off the GP. Consequently, we did not test models that used a constant trend but no GP.; and, finally, 5 different limb darkening models, including two novel parameterizations. All told we attempted to fit 4320 unique light curve/forward model combinations. To ensure our final likelihood values robustly reflect the true global minimum, we repeated each fit 3 times with 3 different initial random seeds. All fits were run on Yale University’s Bouchet high-performance computing cluster.
The remainder of this section is dedicated to a description of each of these models, while the results of the point-estimate fits are presented in Sec. IV.
III.2 Reduction Pipelines and Light Curve Generation
@@ III.2.2 ExoTiC-JEDI
katahdin only includes functionality to process the equivalent of Stage 1 of jwst, meaning it includes no new algorithms for spectral extraction. In an effort to maintain true independence from the other pipelines, however, we opted not to use either jwst, Eureka!, or ExoTiC-JEDI for the equivalent of Stage 3 processing. Instead, we used the following simple routine: looking at just exposure 1, which does not contain any of Kepler-167 e’s transit, we flagged all pixels that fell above various signal-to-noise thresholds and summed them column-wise. This was justified by the minimal background signal left following the GP $1/f$ correction. Our 6 katahdin light curves correspond to using SNR thresholds 3, 5, 8, 10, 15, and 25, where as ordered each light curve relied on fewer and fewer pixels. As discussed in Sec. IV.4, however, we only used the latter four of these in our final nested sampling fits.
@@
III.2.4 Binning/Cleaning Procedure
@@
III.2.5 Pipeline Comparisons
Here we take the opportunity to discuss the white light curves themselves absent any reference to their influence on our fit results. Fig. 1 shows all 18 reductions plotted in full in their 5-minute cadence forms, each normalized by the median of flux in exposure 1. Fig. 2 shows a zoomed-in view of just exposure 1 after dividing out the best-fit quadratic trend from each time series. Two immediate features jump out here: first, light curves that were produced by the same pipeline tend to group together, regardless of their individual pipeline settings. Secondly, all three pipelines tend to agree with one another. Fig. 3 validates both of these conclusions by plotting the correlations between white light curves derived from the same pipeline vs. those from two different pipelines: both cases are positively correlated, but the intra-pipeline distributions are tighter.
@@ III.2.5 Pipeline Comparisons
III.3 Limb Darkening Models
@@ III.3.1 Uninformative Quadratic
III.3.2 Uninformative Cubic
Motivated by the shortcomings of quadratic limb darkening, we also included a more flexible model in our grid: cubic polynomials. Cubics present several practical challenges: for one, one of the most commonly used transit models, Mandel and Agol (2002), is explicitly built around quadratics. This is thankfully circumvented by using a modeling package such as squishyplanet that implements the arbitrary-order limb darkening framework presented in Agol et al. (2020).
@@ III.3.2 Uninformative Cubic
III.3.3 Physically Informed Limb Darkening
@@
To do this, we start by downloading a subset of the pre-computed stellar atmosphere models that underpin exotic-ld (Grant and Wakeford, 2024), an open-source package for computing theoretical limb darkening profiles. Each file corresponds to a star with a certain $T_{\mathrm{eff}}$, log $g$, and metallicity; each file contains several spectra of that star corresponding to different distances from the projected center. We also download the appropriate instrument response curve for our specific filter/grating combination from exotic-ld.
@@
One may fairly ask in all of this why we do not simply fix the limb darkening coefficients before a fit as is occasionally done elsewhere in the literature (e.g. Alam et al., 2022). Doing so, however, runs the risk that a slight misparameterization could irreparably inject or suppress an oblateness signal (and/or alter a planet’s spectrum, Mercier et al., 2025). Further, even if one were to fully trust a given stellar atmosphere model (which would be unwise, given the evidence that they do not perfectly represent observed empirical profiles Espinoza and Jordán, 2015; Patel and Espinoza, 2022), every star comes with uncertainties on its physical characteristics as well.
@@
We have assembled this routine into an open-source python/JAX package we call asterias, after the Latin genus of the common star fish, and hope that the community finds it useful for modeling the limbs of stars. We emphasize that asterias is built on the work of Grant and Wakeford (2024) and their efforts to compile stellar atmosphere model outputs into standardized file types, and that any efforts that rely on asterias should also cite Grant and Wakeford (2024).